# FAM-HRI: Foundation-Model Assisted Multi-Modal Human-Robot Interaction Combining Gaze and Speech

### v1

@@ Abstract
Effective Human-Robot Interaction (HRI) is crucial for enhancing accessibility and usability in real-world robotics applications. However, existing solutions often rely on gestures or language commands, making interaction inefficient and ambiguous, particularly for users with physical impairments. In this paper, we introduce FAM-HRI, an efficient multi-modal framework for human-robot interaction that integrates language and gaze inputs via foundation models. By leveraging lightweight Meta ARIA glasses, our system captures real-time multi-modal signals and utilizes large language models (LLMs) to fuse user intention with scene context, enabling intuitive and precise robot manipulation. Our method accurately determines gaze fixation time interval, reducing noise caused by the gaze dynamic nature. Experimental evaluations demonstrate that FAM-HRI achieves a high success rate in task execution while maintaining a low interaction time, providing a practical solution for individuals with limited physical mobility or motor impairments.
@@
[…]
figureVoice-gaze fusion HRI enables efficient, intuitive interaction without memorization, ideal for physically impaired users.
Supplementary Material
The supplementary materials and video can be found at: https://github.com/laiyuzhi/FAM-HRI
@@ I INTRODUCTION
Individuals with limited mobility or motor impairments, such as neuromuscular disorders, often struggle with daily tasks. These tasks can be automated using robotic manipulators, enhancing independence and improving quality of life [1, 2, 3]. However, improving life through automation requires effective and accessible Human-Robot Interaction (HRI), which remains a challenge. Ensuring smooth interaction is particularly difficult in systems that rely on verbal commands [4]. Accurately processing verbal descriptions becomes problematic when identifying unfamiliar objects or distinguishing between similar items. Existing voice-controlled assistants, such as Amazon Alexa or Tmall Genie, remain insufficient in bridging this gap for embodied HRI.
Gaze, as a natural interaction modality, has great potential in HRI for individuals with severe physical limitations [5, 6, 7, 8], as shown in Fig. FAM-HRI: Foundation-Model Assisted Multi-Modal Human-Robot Interaction Combining Gaze and Speech. However, existing methods often rely on external or bulky devices, restricting mobility and reducing the accuracy of HRI [5, 7, 8]. Some early gaze-based interfaces [6, 9] enable robot control at the cost of increased interaction time and complexity [10]. Beyond hardware constraints, the nature of gaze detection itself adds challenges [11]. During natural gaze interactions, the gaze does not always remain fixed on the target object but instead briefly lingers around the area [12]. This dynamic nature increases noise in target selection process, reducing gaze stability in daily use. These challenges emphasize the need for accurate, efficient, and lightweight systems to enhance assistive robotics.
To address the challenges, we introduce a Foundation-model Assisted Multi-modal HRI framework, FAM-HRI, that enables users to convey their intentions to robots. By wearing lightweight glasses from the Meta ARIA Research Kit [13], the system can simultaneously collect gaze and voice signals from users. By integrating these modalities, the robot accurately selects target objects and executes intended actions based on environmental context, reducing ambiguity in language-based systems, especially for users with limited mobility. A command-processing LLM analyzes user input to identify the optimal gaze period, extracting high-quality signals while minimizing noise from gaze dynamics. A motion-planning LLM then generates the robot’s action sequences, leveraging advanced reasoning to handle complex contexts more effectively than traditional rule-based systems. Our main contributions are summarized below:
  • To the best of our knowledge, FAM-HRI is the one of the first frameworks to integrate voice and gaze with a lightweight wearable device for human-robot collaboration. By leveraging ARIA glasses, our approach facilitates natural and intuitive communication, enhancing accessibility and ease of use. We assess its effectiveness through a comprehensive user experience study.
@@
  • Our system optimizes gaze fixation timing and robot planning by integrating language, gaze, and environmental context. This ensures precise user intent alignment with robot actions, enhancing task execution.
  • FAM-HRI was benchmarked against state-of-the-art HRI methods, achieving higher interaction success rates and faster input processing. To support the community, we will release our system design, algorithms, and solutions as open-source at github.com/laiyuzhi/FAM-HRI.
II Related Work
@@
Among alternative modalities, gaze offers a powerful interaction method for individuals with disabilities. Most existing gaze-based assistive devices rely on graphical user interface (GUI) systems, which enable users to control assistive devices through visual interfaces. Some researchers employ virtual buttons, allowing users to operate electric wheelchairs [9], robotic arms [23], or mobile robots [24]. However, GUI-based systems divert user attention from the external environment, limiting their field of view. An optimal system should enable unrestricted gaze-based interaction by accurately estimating user attention [10]. In [25], researchers introduced a gaze-following model that predicts gaze direction and visual attention using RGB images from a robot’s perspective combined with YOLO-based object detection. Similarly, [26] applies transfer learning to track gaze within predefined regions, allowing the system to determine attention states. This enhances the ability of the robot to interpret user intentions, improving efficiency and naturalness in human-robot collaboration, particularly in industrial settings. However, these approaches also require users to fixate on targets continuously, which feels unnatural compared to typical communication, where the gaze naturally shifts.
Existing HRI approaches remain inadequate for general use, particularly for individuals with difficulties in language commands or physical movement. Integrating LLMs with strong reasoning capabilities offers potential for enhancing multi-modal data processing across various HRI applications, including social robots [18, 19] and robot navigation [27]. Recently, researchers introduced LaMI [28] to explore how LLMs can improve multi-modal HRI by reasoning over diverse inputs. However, this approach primarily focuses on backend reasoning and overlooks frontend measurement uncertainties and other external factors affecting perception and decision-making in real-world interactions.
@@ III Problem Definition
Our proposed system integrates environmental perception from both the human and the robot points of view to facilitate effective collaboration. The human view is mainly used for intention fusion, capturing user-centric cues, while the robot view focuses on manipulation, allowing precise interaction with objects. To structure multi-modal signals for foundation model reasoning, we need to define sets of geometric frames and control parameters.
@@ III-A Definitions of Frames and Transformations
To ensure consistency, we define the reference frames used in this paper. The robot base frame ()$\prescript{r}{}{(\cdot)}$ serves as the primary reference for the robot, while the robot camera frame ()$\prescript{c}{}{(\cdot)}$ is used for manipulation. The ARIA glasses camera frame (()$\prescript{gc}{}{(\cdot)}$) provides a first-person user-centric perspective. The ARIA pupil frame (()$\prescript{gp}{}{(\cdot)}$), centered at the midpoint of the glasses’ bridge, serves as the reference for gaze vector estimation. The transformation $T\in\mathbb{SE}(3)$ denotes a rigid body transformation in the special Euclidean group. Specifically, the transformation matrix ${}^{gc}T_{gp}$ maps points from the ARIA pupil frame to the ARIA camera frame, with parameters provided by the manufacturer111facebookresearch.github.io/projectaria_tools/.
@@ III-B Formulation of Human View Representation
To formally characterize the human view inputs of ARIA glasses, we define the human input state vector as $\mathcal{Z}_{H}=(\mathcal{S},\mathcal{U},\mathcal{G})$. Here, $\mathcal{S}\in\Sigma^{*}$ denotes the transcribed speech sequence. The audio is recorded at a 48 kHz sampling rate using the built-in microphone of the ARIA glasses and converted into text via Fast Whisper [29], with each word timestamped. Six embedded microphones improve speech clarity by suppressing background noise and improving transcription accuracy even in noisy environments. Additional details are provided in the supplementary material. $\mathcal{U}\in\mathbb{R}^{H_{1}\times W_{1}\times(t_{n}-t_{1})}$ represents the RGB image sequence captured over the interval $(t_{n}-t_{1})$ by the user-centric camera (Resolution: $1408\times 1408$, 10 FPS). $\mathcal{G}\in\mathbb{R}^{H_{2}\times W_{2}\times(t_{n}-t_{1})}$ corresponds to the grayscale eye-tracking image sequence (Resolution: $320\times 240$ per eye, 10 FPS), independently capturing the left and right eyes for gaze estimation.
The human-view inputs are utilized to infer the referred object, represented as $\overline{\mathcal{Z}_{g}}=(\overline{\mathbf{p}_{g}},\ \overline{\beta_{g}},\ \overline{M_{g}}),$ which encodes its spatial and visual properties. Here, $\overline{\mathbf{p}_{g}}\in\mathbb{R}^{2}$ denotes the 2D position of the referred object, $\overline{\beta_{g}}\in\mathbb{R}^{4}$ defines its bounding box, and $\overline{M_{g}}\in\mathbb{R}^{H_{1}\times W_{1}}$ represents its segmentation mask.
@@ III-C Formulation of Robot View Representation
Figure 2: Human view intention fusion and multi-view intention alignment
@@ III-D Formulation of Control System
The control system is responsible for generating and executing actions based on the intention of the human and robot views. The system is manipulated with a predefined set of parameterized action primitives $\mathcal{A}(\Theta)=\left\{a_{i}(\theta_{i})\mid i\in\mathbb{R}\right\}$ and follows a parameterized planning policy to ensure adaptive and context-aware manipulation. In $\mathcal{A}(\Theta)$ each primitive action $a$ is parameterized by 2D position, object category, translation distance, and rotation angle. These action primitives form low-level and high-level control commands, enabling the robot to execute pick-up, placing, pushing, moving, and rotating actions in a structured manner.
@@
We propose a framework for multi-modal HRI combined language and gaze via foundation models. Our proposed system diagram is shown in Fig. 1.
@@ IV Methodology
Our system enables multi-modal HRI by integrating gaze, language, and visual perception across both human and robot viewpoints to achieve accurate intention inference and action execution. The methodology is structured into three key components: Human View Intention Fusion, Multi-View Intention Alignment, and Planning Policy Generation.
@@ IV-A1 GPT-Agent for Human View Command Processing
where $\mathcal{L}$ defines the target properties, with possible values “object” or “position”. The variable $c_{target}$ specifies the target object category in the command, later utilized for scene observation in both human and robot views. If the referenced is a general category, the output is “stuff”; if it specifies a position without an object, it is labeled “position”; otherwise, it corresponds to a specific object category. $\Delta t$ indicates the time period of the word spoken by the user determined by Fast Whisper [29], during which the user is most likely fixating on the target object or location. For instance, given the command: “please put the apple there on the table”. The system interprets it as: $\mathbf{O}_{1}=\left\{(object,\ position),\ (apple,\ table),\ (\Delta t_{apple},\ \Delta t_{there})\right\}$. This structured output aligns gaze fixation with spoken language to enhance intention inference.
@@ IV-A2 Scene Observation in Human View
We define a scene observation set in human view $\mathcal{Z}_{g}$ to describe target object categories spatial and visual properties as $\mathcal{Z}_{g}=\left\{c_{target},\ \{\mathbf{p}_{g_{i}},\ \beta_{g_{i}},\ M_{g_{i}}\}_{i=1}^{N}\right\}$. In this set $\mathbf{p}_{g_{i}}\in\mathbb{R}^{2}$, $\beta_{g_{i}}\in\mathbb{R}^{4}$ and $M_{g_{i}}\in\mathbb{R}^{H_{1}\times W_{1}}$ are the 2D position, bbox and mask of the target object $i$ and $c_{target}$ is the category for target object from the first GPT-Agent. $\mathbf{p}_{g_{i}}$ is defined as the midpoint of the $\beta_{g_{i}}$. A total of $N$ objects of this category are observed. In our proposed approach, Grounding Dino SAM2 [30, 31] performs object detection, classification, and segmentation with prompt $c_{target}$ from Eq. 2.
@@
For 3D gaze estimation, the ARIA glasses incorporate two eye-tracking cameras. We employ the SocialEye model222github.com/facebookresearch/projectARIA_eyetracking to predict the 3D gaze vector ${}^{gp_{t_{i}}}\mathbf{P}_{{t_{i}}}\in\mathbb{R}^{3}$ at time $t_{i}$ in the pupil frame of the glasses. To reconstruct the 3D gaze vector within $\Delta t$ at time $t_{i+n}$, donated as ${}^{gc_{t_{i+n}}}\mathbf{P}_{{t_{i+n}}}$, onto the frame ()$\prescript{gc}{}{(\cdot)}$ at time $t_{i}$, denoted as ${}^{gc_{t_{i}}}\mathbf{P}_{{t_{i+n}}}\in\mathbb{R}^{3}$, and then project them onto image plane ${}^{{t_{i}}}\mathcal{U}$ at $t_{i}$ as gaze point ${}^{gc_{t_{i}}}\mathbf{p}^{gaze}_{{t_{i+n}}}\in\mathbb{R}^{2}$ we use the transformations:

### v2

@@ Abstract
Effective Human-Robot Interaction (HRI) is crucial for enhancing accessibility and usability in real-world robotics applications. However, existing solutions often rely on gestures-only or language-only commands, making interaction inefficient and ambiguous, particularly for users with physical impairments. In this paper, we introduce FAM-HRI, an efficient multi-modal framework for human-robot interaction that integrates language and gaze inputs via foundation models. By leveraging lightweight Meta ARIA glasses, our system captures real-time multi-modal signals and utilizes large language models (LLMs) to fuse user intention with scene context, enabling intuitive and precise robot manipulation. Our method accurately determines gaze fixation time interval, reducing noise caused by the gaze dynamic nature. Experimental evaluations demonstrate that FAM-HRI achieves a high success rate in task execution while maintaining a low interaction time, providing a practical solution for individuals with limited physical mobility or motor impairments. To support the community, we will release our system design, algorithms, and solutions as open-source after paper acceptance.
Note to Practitioners
In the field of assistive and service robotics, enabling users with physical disabilities to communicate intentions naturally remains a significant challenge. Existing uni-modal approaches, such as voice-only or gaze-only interaction, often struggle with ambiguity and environmental interference, limiting their reliability in real-world scenarios. Current multi-modal systems typically process gaze and speech separately, making it difficult to align user’s visual attention with spoken commands accurately. Additionally, many systems rely on bulky hardware or require users to maintain prolonged focus, reducing practicality and comfort. To address these issues, this paper proposes a foundation-model-assisted multi-modal human-robot interaction (FAM-HRI) framework that combines lightweight AR glasses with LLMs to fuse gaze and speech inputs in real time. The system automatically filters noisy gaze signals, aligns them with spoken commands, and generates precise robot actions. Experiments demonstrate that this approach achieved task success rates exceeding 94 % across multiple scenarios and outperformed all baseline methods in interaction efficiency, with the shortest interaction times in comparative evaluations. This method is well-suited for applications such as household assistance, and collaborative tasks where hands-free interaction is desirable. In the future, we plan to optimize computing performance and explore strategies to resolve conflicts between gaze and speech inputs, further enhancing usability and adaptability in diverse environments.
Index Terms:
Robot vision system, Robot control, Perception, Cooperative systems.
@@
Figure 1: Voice-gaze fusion HRI enables efficient, intuitive interaction without memorization, ideal for physically impaired users.
@@ I Introduction
In recent years, assistive robotics has attracted increasing attention due to its potential to enhance the autonomy and quality of life of individuals with physical disabilities or neuromuscular impairments. Tasks that most people perform effortlessly—like reaching for a cup or opening a drawer—can be daunting for users with limited motor control. Robotic manipulators offer a promising path to automate these actions [1, 2, 3]. However, realizing this vision hinges on effective and inclusive Human-Robot Interaction (HRI). Traditional interfaces, such as joysticks or touchscreen panels, demand dexterous control and are thus often inaccessible. Voice-controlled systems are more inclusive, yet remain insufficient in accurately interpreting users’ intentions, particularly when referring to unfamiliar objects or choosing among similar items [4, 5]. Additionally popular voice-driven assistants like Amazon Alexa or Tmall Genie lack embodied grounding and often fail in real-world complex tasks. Thus, there is a pressing need for more intuitive, context-aware interaction paradigms that can bridge the gap between human intent and robot execution.
Recent works [6, 7, 8, 9] have explored alternative interaction modalities such as eye gaze. Gaze is a natural, low-effort cue that can reveal intent without requiring physical movement. Early gaze-driven systems have demonstrated feasibility in robot control [7, 10], yet they rely on bulky external trackers or head-mounted cameras that impede mobility and resist integration into daily environments [6, 8, 9]. Moreover, combining gaze with traditional Graphical User Interface (GUI) elements increases interaction time and complexity [11]. While some approaches attempt to fuse gaze with speech, they typically process the two modalities independently, missing critical cross-modal correlations. As a result, most systems fail to deliver both the precision and ease-of-use needed for everyday assistive HRI, especially in cluttered and dynamic environments of real-world homes.
Despite its potential, leveraging gaze and speech in HRI introduces several fundamental challenges: (1) Natural gaze behavior is dynamic and noisy: the user’s eyes do not remain fixed but instead shift subtly around the target area [12], introducing temporal noise and uncertainty in object selection. Gaze patterns may reflect intention only briefly, complicating efforts to extract meaningful cues. (2) Verbal commands are often ambiguous—especially when users must describe unknown or visually similar items without using precise labels [5]. (3) Existing systems generally require expensive or immobile hardware setups, which hinder deployment in daily living environments [6, 8]. (4) Few methods model the temporal alignment between gaze and speech. In real-world scenarios, a user may look at an object moments before or after issuing a verbal command. Without a principled way to synchronize and interpret these signals, errors propagate and degrade the robot’s performance. These challenges underscore the need for lightweight, intelligent, and modality-coordinated interaction frameworks.
To overcome these limitations, we propose FAM-HRI—a Foundation-model-Assisted Multi-modal Human-Robot Interaction framework that enables intuitive communication for users with limited mobility. Our system leverages lightweight wearable glasses from the Meta ARIA Research Kit [13], which can simultaneously capture egocentric video, gaze signals, and voice commands. A foundation language model agent processes the spoken command and selects the most semantically aligned gaze period, filtering out irrelevant or noisy frames caused by microsaccades or transient fixations. Another large model agent performs high-level motion planning based on multi-modal context and environmental cues. As a result, users can simply glance at their intended object and speak naturally—the system then robustly identifies the target and commands the robot to act. This joint modeling of gaze and speech enables more precise, adaptive, and accessible HRI, particularly in complex scenes. By reducing ambiguity and minimizing user effort, FAM-HRI bridges the gap between intent and robotic autonomy in assistive settings.
Our main contributions are summarized below:
  • To the best of our knowledge, FAM-HRI is one of the first frameworks to integrate simultaneous voice and gaze with a lightweight wearable device for efficient human-robot collaboration. Unlike prior works that treats gaze and speech as separate inputs [11, 14], our proposed unified multi-modal HRI framework improves robustness and precision by aligning gaze and speech commands, enabling reliable disambiguation in complex environments.
@@
  • Our system optimizes gaze fixation timing and robot planning by integrating language, gaze, and environmental context. This ensures precise user intent alignment with robot actions, enhancing task execution. The experimental results illustrate that users can complete complex tasks with a high success rate with only brief instruction.
  • Compared with the state-of-the-art HRI methods, FAM-HRI achieves higher interaction success rates and faster input processing. By leveraging ARIA glasses, our approach facilitates natural and intuitive communication, enhancing accessibility and ease of use. We assess its effectiveness through a comprehensive user experience study.
II Related Works
@@
Among alternative modalities, gaze offers a powerful interaction method for individuals with disabilities. Most existing gaze-based assistive devices rely on graphical user interface (GUI) systems, which enable users to control assistive devices through visual interfaces. Some researchers employ virtual buttons, allowing users to operate electric wheelchairs [10], robotic arms [23], or mobile robots [24]. However, GUI-based systems divert user attention from the external environment, limiting their field of view. An optimal system should enable unrestricted gaze-based interaction by accurately estimating user attention [11]. Recent work such as [14] focuses on robust gaze-based intention prediction in real-world scenarios. Their method applies OPTICS clustering to filter out redundant gaze points, enabling robust intention prediction.However, this approach presents several limitations. First, it is not integrated with an actual robot control system, and thus lacks validation in closed-loop interaction. Second, it does not consider the alignment between gaze and language input; users can only select target objects through gaze, limiting natural multi-modal expression. In [25], researchers introduced a gaze-following model that predicts gaze direction and visual attention using RGB images from a robot’s perspective combined with YOLO-based object detection. Similarly, [26] applies transfer learning to track gaze within predefined regions, allowing the system to determine attention states. This enhances the ability of the robot to interpret user intentions, improving efficiency and naturalness in human-robot collaboration, particularly in industrial settings. In addition, some researches in [27] proposed a human-in-the-loop multi-modal fusion method that dynamically adjusts the weights of gaze, gesture, and speech modalities based on user feedback. This adaptive mechanism improves the system’s robustness by resolving conflicting inputs during interaction. However, the approach is limited to a fixed set of predefined object categories and supports only simple, hard-coded robot actions. As a result, it lacks the flexibility to engage in natural, open-ended interaction and cannot support more complex or extended task sequences. However, these approaches also require users to fixate on targets continuously, which feels unnatural compared to typical communication, where the gaze naturally shifts.
Existing HRI approaches remain inadequate for general use, particularly for individuals with difficulties in language commands or physical movement. Integrating LLMs with strong reasoning capabilities offers potential for enhancing multi-modal data processing across various HRI applications, including social robots [5, 19] and robot navigation [28]. Recently, researchers introduced LaMI [29] to explore how LLMs can improve multi-modal HRI by reasoning over diverse inputs. While LaMI provides a strong backend reasoning capability, it primarily focuses on symbolic-level interpretation based on human pose estimation, dialogue history and scene detection. This approach overlooks frontend measurement uncertainties and other external factors affecting perception and decision-making in real-world interactions. For example, LaMI infers the user’s attention direction based on head orientation to determine whether the user is referring to the robot or the workspace, rather than determining the specific object to which the user is visually attending.
@@
Our proposed system integrates environmental perception from both the human and the robot points of view to facilitate effective collaboration. The human view is mainly used for intention fusion, capturing user-centric cues, while the robot view focuses on manipulation, allowing precise interaction with objects.
@@ III Problem Definition
Our proposed system integrates environmental perception from both human and robot views to facilitate effective collaboration. The human view is mainly used for intention fusion, capturing user-centric cues, while the robot view focuses on manipulation, allowing precise interaction with objects. To structure multi-modal signals for foundation model reasoning, we formalize sets of geometric frames and control parameters. For comprehensive definitions for the mathematical notations used throughout our paper, please refer to Appendix C.
@@ III-A Definitions of Frames and Transformations
To ensure consistency, we define the reference frames used in this paper. The robot base frame ()$\prescript{r}{}{(\cdot)}$ serves as the primary reference for the robot, while the robot camera frame ()$\prescript{c}{}{(\cdot)}$ is used for manipulation. The ARIA glasses camera frame ()$\prescript{gc}{}{(\cdot)}$ provides a first-person user-centric perspective. The ARIA pupil frame ()$\prescript{gp}{}{(\cdot)}$, centered at the midpoint of the glasses’ bridge, serves as the reference for gaze vector estimation. The transformation $T\in\mathbb{SE}(3)$ denotes a rigid body transformation in the special Euclidean group. Specifically, the transformation matrix ${}^{gc}T_{gp}$ maps points from the ARIA pupil frame to the ARIA camera frame, with parameters provided by the manufacturer111facebookresearch.github.io/projectaria_tools/.
@@ III-B Formulation of Human View Representation
To formally characterize the human view inputs of ARIA glasses, we define the human input state vector as $\mathcal{Z}_{H}=(\mathcal{S},\mathcal{U},\mathcal{G})$. Here, $\mathcal{S}\in\Sigma^{*}$ denotes the transcribed speech sequence. The audio is recorded at a 48 kHz sampling rate using the built-in microphone of the ARIA glasses and is converted into text via Fast Whisper [30], with each word timestamped. Six embedded microphones improve speech clarity by suppressing background noise and improving transcription accuracy even in noisy environments. $\mathcal{U}\in\mathbb{R}^{H_{1}\times W_{1}\times(t_{n}-t_{1})}$ represents the RGB image sequence captured over the interval $(t_{n}-t_{1})$ by the user-centric camera (Resolution: $1408\times 1408$, 10 FPS). $\mathcal{G}\in\mathbb{R}^{H_{2}\times W_{2}\times(t_{n}-t_{1})}$ corresponds to the grayscale eye-tracking image sequence (Resolution: $320\times 240$ per eye, 10 FPS), independently capturing the left and right eyes for gaze estimation.
The human-view inputs are utilized to infer the referred object, represented as $\overline{\mathcal{Z}_{g}}=(\overline{\mathbf{p}_{g}},\ \overline{\beta_{g}},\ \overline{M_{g}}),$ which encodes its spatial and visual properties. Here, $\overline{\mathbf{p}_{g}}\in\mathbb{R}^{2}$ denotes the 2D position of the referred object, $\overline{\beta_{g}}\in\mathbb{R}^{4}$ defines its bounding box, and $\overline{M_{g}}\in\mathbb{R}^{H_{1}\times W_{1}}$ represents its segmentation mask. $\overline{\mathbf{p}_{g}}$ is defined as the midpoint of the $\overline{\beta_{g}}$.
@@ III-D Formulation of Control System
The control system is responsible for generating and executing actions based on the intention of the human and robot views. The system is manipulated with a predefined set of parameterized action primitives $\mathcal{A}(\Theta)=\left\{a_{i}(\theta_{i})\mid i\in\mathbb{R}\right\}$ and follows a parameterized planning policy to ensure adaptive and context-aware manipulation. In $\mathcal{A}(\Theta)$, each primitive action $a$ is parameterized by 2D position, object category, translation distance, and rotation angle. These action primitives form low-level and high-level control commands, enabling the robot to execute pick-up, placing, pushing, moving, and rotating actions in a structured manner.
@@ IV Methodology
Our system enables multi-modal HRI by integrating gaze, language, and visual perception across both human and robot viewpoints to achieve accurate intention inference and action execution. The methodology is structured into three key components: Human View Intention Fusion, Multi-View Intention Alignment, and Planning Policy Generation. Our proposed system diagram is shown in Fig. 2.
@@ IV-A1 GPT-Agent for Human View Command Processing
where $\mathcal{L}$ defines the target properties, with possible values “object” or “position”. The variable $c_{target}$ specifies the target object category in the command, later utilized for scene observation in both human and robot views. If the referenced target is a general category, the output is “stuff”; if it specifies a position without an object, it is labeled “position”; otherwise, it corresponds to a specific object category. $\Delta t$ indicates the time period of the word spoken by the user determined by Fast Whisper [30], during which the user is most likely fixating on the target object or location. For instance, given the command: “please put the apple there on the table”. The system interprets it as: $\mathbf{O}_{1}=\left\{(object,\ position),\ (apple,\ table),\ (\Delta t_{apple},\ \Delta t_{there})\right\}$. This structured output aligns gaze fixation with spoken language to enhance intention inference.
@@ IV-A2 Scene Observation in Human View
Figure 3: Schematic of 3D gaze estimation and reconstruction.
We define a scene observation set in human view $\mathcal{Z}_{g}$ to describe target object categories spatial and visual properties as $\mathcal{Z}_{g}=\left\{c_{target},\ \{\mathbf{p}_{g_{i}},\ \beta_{g_{i}},\ M_{g_{i}}\}_{i=1}^{N}\right\}$. In this set $\mathbf{p}_{g_{i}}\in\mathbb{R}^{2}$, $\beta_{g_{i}}\in\mathbb{R}^{4}$ and $M_{g_{i}}\in\mathbb{R}^{H_{1}\times W_{1}}$ are the 2D position, bbox and mask of the target object $i$ and $c_{target}$ is the category for target object from the first GPT-Agent. $\mathbf{p}_{g_{i}}$ is defined as the midpoint of the $\beta_{g_{i}}$. A total of $N$ objects of this category are observed. In our proposed approach, Grounding DINO [31] and SAM 2 [32] perform object detection, classification, and segmentation with prompt $c_{target}$ from Eq. 2.
@@
For 3D gaze estimation, the ARIA glasses incorporate two eye-tracking cameras. We employ the SocialEye model222github.com/facebookresearch/projectARIA_eyetracking to predict the 3D gaze vector ${}^{gp_{t_{i}}}\mathbf{P}_{{t_{i}}}\in\mathbb{R}^{3}$ at time $t_{i}$ in the pupil frame of the glasses. As shown in Fig. 3, to reconstruct the 3D gaze vector within $\Delta t$ at time $t_{i+n}$, donated as ${}^{gc_{t_{i+n}}}\mathbf{P}_{{t_{i+n}}}$, onto the frame ()$\prescript{gc}{}{(\cdot)}$ at time $t_{i}$, denoted as ${}^{gc_{t_{i}}}\mathbf{P}_{{t_{i+n}}}\in\mathbb{R}^{3}$, and then project them onto image plane ${}^{{t_{i}}}\mathcal{U}$ at $t_{i}$ as gaze point ${}^{gc_{t_{i}}}\mathbf{p}^{gaze}_{{t_{i+n}}}\in\mathbb{R}^{2}$ we use the transformations: